# Concept-Aware LoRA for Domain-Aligned Segmentation Dataset Generation

### v1

@@ 4 Experiments
This section presents extensive experiments on urban-scene segmentation in both in-domain and domain generalization (DG) settings. Section 4.1 describes the experimental setup and implementation details. Then, Section 4.2 presents extensive urban-scene segmentation experiments across in-domain and DG settings. Finally, Section 4.3 provides an in-depth analysis of the Concept-Aware LoRA (CA-LoRA), including a comprehensive ablation study.
Figure 5: Qualitative comparison of image-label pairs between DatasetDM, Original LoRA, and our CA-LoRAs in a few-shot setting (Cityscapes, 0.3%). The pretrained model misaligns with the viewpoint and style of the target domain, while the original LoRA memorizes training examples. In contrast, CA-LoRA selectively learns either style or viewpoint concepts from the source dataset.
@@ 4.1 Experimental Setup
Datasets
For training, we use Cityscapes [9] as the source dataset for all urban-scene segmentation experiments, including in-domain and DG settings. Regarding the experiments for the in-domain few-shot setting, we only utilize a subset of Cityscapes images for the few-shot samples. For the evaluation, we test on Cityscapes (CS) validation set for in-domain, and ACDC [45], Dark Zurich (DZ) [44], BDD100K (BDD) [57], and Mapillary Vistas (MV) [33] for DG settings. Notably, ACDC and DZ are constructed with adverse weather conditions. We conduct additional experiments on Pascal-VOC [12] for general segmentation, with results provided in Supplementary Section D.
@@
For baseline model, we train Mask2Former [6] on subsets of the Cityscapes [9] dataset at various fractions (0.3%, 1%, 3%, 10%, and 100%), which we denote as ‘Baseline’. Then, we generate 500 image-label pairs for all few-shot settings and use them as an additional dataset to fine-tune the baseline model, where we generate 3000 pairs for the fully-supervised setting. To prevent overfitting to the generated data, we mix real (e.g., samples from the Cityscapes dataset) and generated samples in equal proportions within each mini-batch. We compare our proposed approach with the segmentation dataset generation approach, DatasetDM [51], which utilizes the pretrained T2I model [41] without any fine-tuning. We include an additional baseline denoted as ‘Baseline (FT)’ that fine-tunes the baseline model exclusively on the same real dataset, ensuring a fair comparison in terms of the number of training iterations.
@@
We improve the DG performance for urban-scene segmentation by building upon existing DG methods, including ColorAug, DAFormer [19] and HRDA [20]. To show the effectiveness of the generated dataset, we train a semantic segmentation model with each DG method on a combination of the 2975 Cityscapes image-label pairs and the 2500 generated image-label pairs (500 images for 5 weather conditions including clear, foggy, night-time, rainy, and snowy) from scratch. We compare our approach with various dataset generation approaches, including DatasetDM [51], DGInStyle [23]. DATUM [4], and InstructPix2Pix [5]. 444 For DATUM [4], we provide an additional image for each weather condition to meet the requirements of the one-shot UDA setting. For InstructPix2Pix [5], we modify the weather conditions of the images using instruction prompts (e.g., “change the weather condition to {…}”) while preserving the original segmentation maps.
Implementation Details
Throughout our experiments, we use Stable Diffusion XL [37] as the pretrained T2I model. For a fair comparison, we reimplement DatasetDM [51] with SDXL as well. We fix the rank of both the original LoRA and CA-LoRA at 64 and set the training iterations to 10k. While our method requires additional fine-tuning of the T2I model, increasing training time compared to DatasetDM, this process (i.e., CA-LoRA fine-tuning) takes only one hour on a single Tesla V100 GPU, incurring minimal computational cost. Thus, its overhead remains marginal compared to the 20-hour training time of the label generator. We explore the selected proportion of CA-LoRA across 1%, 2%, 3%, 5%, and 10%, since 0% and 100% represent DatasetDM and the original LoRA, respectively. The diffusion timestep $t$ for identifying the desired concept is set to 81 out of 1000 timesteps. The results of our hyper-parameter search, along with details such as the label generator architecture, are provided in Section G.
@@ 4.2 Main Results on the Segmentation Benchmarks
This section demonstrates the superiority of CA-LoRA in improving urban-scene segmentation performance through both quantitative and qualitative results.
@@
Table 1 shows that the proposed CA-LoRA consistently outperforms all other methods across various data ratios. Specifically, the proposed method improves mIoU by 2.30 for the 0.3% fraction of the Cityscapes, significantly surpassing DatasetDM, which achieves only a 0.99 mIoU increase. This demonstrates the efficiency of CA-LoRA, particularly in low-data regimes. Furthermore, our method improves the fully-supervised performance (i.e., 100% data ratio) by 1.34 mIoU, further enhancing strong Mask2Former performance. The consistent gains across different data ratios highlight the robustness of CA-LoRA, making it highly effective for urban-scene segmentation in both few-shot and fully supervised settings.
@@
As shown in Table 2, the proposed method consistently outperforms all other segmentation dataset generation methods across multiple DG methods. Specifically, Viewpoint CA-LoRA effectively learns only the viewpoint from the source dataset (Cityscapes) while maintaining the ability to generate diverse styles from the pretrained T2I model. As a result, our method significantly improves generalization performance, particularly on challenging datasets such as ACDC and Dark Zurich, where conditions such as adverse weather play a critical role. We emphasize that DGInStyle and InstructPix2Pix only change the styles of given images while keeping fixed label maps, which introduces limited manipulation of the scene. Since DAFormer and HRDA already employ strong image augmentation techniques [19, 20], the additional simplistic image augmentations from DGInStyle and InstructPix2Pix are largely redundant, as shown in their performance. In contrast, the improvement of the proposed approach is notable not only in comparison to the simple baseline (ColorAug) but also in its effectiveness on the advanced DG methods (DAFormer and HRDA) since our generated datasets contain semantically diverse content.
Metric Desired Concept Proportion of T2I Model Layers for Concept-Aware Fine-tuning
0% (Pretrained) 1% 2% 3% 5% 10% 100% (Original LoRA)
CMMD Style 5.063 1.618 1.420 1.021 1.105 0.686 0.644
Viewpoint 2.650 2.313 1.580 1.733 1.476
CLIP Score Style 25.72 21.44 19.92 20.74 19.72 19.86 22.66
Viewpoint 25.03 24.69 25.88 25.52 24.76
Image-Label Alignment (mIoU) Style 25.18 37.01 39.37 40.94 47.84 48.86 60.10
Viewpoint 33.27 29.42 31.44 34.94 38.28
Table 3: Evaluating generated datasets according to the desired concepts and the proportion of T2I model layers for concept-aware finetuning in a few-shot setting (Cityscapes, 0.3%). The generated datasets are measured by image domain alignment (CMMD $\downarrow$), fidelity of the weather-augmented prompts (CLIP-Score $\uparrow$), and image-label alignment (mIoU $\uparrow$).
Additional Generated Dataset In-domain (0.3%) Domain Generalization (ColorAug)
Cityscapes ACDC DZ BDD MV Average
- 41.83 53.12 25.69 53.00 59.81 47.91
Pretrained (DatasetDM) 42.82 (+0.99) 53.80 27.70 53.54 60.75 48.95 (+1.04)
Original LoRA 42.97 (+1.14) 54.25 28.42 54.34 61.42 49.61 (+1.70)
Style CA-LoRA 44.13 (+2.30) 52.55 26.42 54.04 61.81 48.71 (+0.80)
Viewpoint CA-LoRA 43.13 (+1.30) 56.07 29.75 54.35 61.40 50.39 (+2.48)
Table 4: Ablation study of CA-LoRA on the in-domain (Cityscapes, 0.3%) and domain generalization (ColorAug, CS $\rightarrow$ ACDC, DZ, BDD, MV) settings (mIoU). Style CA-LoRA demonstrates significant improvements in the in-domain setting and Viewpoint CA-LoRA outperforms in the domain generalization setting.
4.3 Analysis
This section presents an in-depth analysis of CA-LoRA finetuned models (style and viewpoint), comparing them to the pretrained base model and the original LoRA finetuned model. We first analyze the effective style adaptation of the Style CA-LoRA, then examine whether Viewpoint CA-LoRA maintains conditional image generation ability. Finally, we show a comprehensive ablation study for in-domain and DG for urban-scene segmentation across the pretrained model, CA-LoRA finetuned models (style, viewpoint), and the original LoRA finetuned model. Further analyses of concept sensitivity and CA-LoRA are provided in Supplementary Sections B and C, respectively. Class-specific generation, including focusing on a specific class and exploring diverse class names, is presented in Supplementary Section F.
@@ 4.3 Analysis
We evaluate the image domain alignment of T2I models, which is crucial for generating an in-domain dataset. To assess the image domain alignment, we adopt CMMD [22] due to the reliability of maximum mean discrepancy in comparing small image sets. Table 3 shows that the pretrained T2I model exhibits a significant domain gap between real and generated images, whereas fine-tuning on Cityscapes effectively reduces this gap. Among CA-LoRAs, Style CA-LoRA achieves competitive image domain alignment with only 10% of the fine-tuned parameters of the original LoRA. While the original LoRA performs best in the CMMD metric, Fig. 1 highlights its memorization issue, which we further analyze in the following ablation study to demonstrate the limitations of the memorized dataset.
@@
Since generating diverse weather conditions (e.g., foggy, rainy) is crucial for improving DG performance, preserving the text adherence ability of the T2I model is essential. We measure the weather-conditional image generation performance by leveraging CLIP-Score [38], which can assess the similarity between the generated images and their input text prompts. Table 3 shows the average CLIP-Score across four diverse weather conditions, based on 100 generated images per condition. The results indicate that the pretrained model achieves a high CLIP score for generating adverse weather conditions, whereas the original LoRA and Style CA-LoRA fail to do so, as they primarily learn styles from the source dataset, which inherently includes its weather condition (e.g., clear-day weather). In contrast, the Viewpoint CA-LoRA effectively preserves the adverse weather conditional generation performance while learning the viewpoint from the source dataset.
@@
To further demonstrate the effectiveness of our method, we analyze image-label alignment. As shown in Fig. 5 and Table 3, our method outperforms baselines both qualitatively and quantitatively by fine-tuning the T2I model with CA-LoRA for domain alignment. This ensures that generative features at inference are similar to those used during label generator training. To quantitatively measure performance with mIoU, we use the predicted label of Mask2Former [6] from the generated image of the T2I model as pseudo ground truth. Details are provided in Supplementary Section E.
Ablation Study
We conduct an ablation study on CA-LoRA for both in-domain (Cityscapes, 0.3%) and DG (CS $\rightarrow$ ACDC, DZ, BDD, MV) settings. As shown in Table 4, the pretrained model and original LoRA show poor performance due to domain misalignment and memorization, respectively. In contrast, our Style CA-LoRA achieves a significant performance improvement of 2.30 by learning in-domain styles while preserving pretrained knowledge, enabling the generation of diverse scenes. Furthermore, as shown in Table 4, the Viewpoint CA-LoRA shows a significant improvement of 2.48 on average for domain generalization. While the original LoRA and Style CA-LoRA show competitive performance improvements on BDD and MV, Viewpoint CA-LoRA significantly outperforms ACDC and DZ, which contain images under challenging weather conditions. These results show the strength of Viewpoint CA-LoRA in synthesizing adverse weather conditions.
@@ 5 Conclusion and Future Work
This paper proposes Concept-Aware LoRA (CA-LoRA), a novel fine-tuning method that learns only the desired concepts (e.g., viewpoint or style) from the training dataset to generate segmentation datasets. By selectively identifying and updating only the weights relevant to the desired concepts, CA-LoRA enables the fine-tuned T2I model to produce well-aligned and informative samples. Moreover, it effectively leverages pretrained knowledge from T2I models to generate datasets, significantly improving segmentation performance across various settings, including in-domain (few-shot and fully-supervised) and domain generalization tasks. While our experiments primarily focus on urban-scene segmentation, additional experiments on PASCAL-VOC [12] demonstrate the general applicability of CA-LoRA (Supplementary Section D). Furthermore, CA-LoRA excels at learning only the desired concepts from training data, even when it contains diverse information, including unnecessary concepts for dataset generation, making it applicable to various knowledge transfer and fine-tuning tasks. As a promising future direction, our approach could be extended to extract specific concepts beyond style and viewpoint, enabling more personalized image generation.
@@ A Appendix
In the supplementary material, we provide additional analyses, experiments, qualitative and quantitative results, and implementation details, including pseudocode, due to space limitations in the main manuscript. Specifically, we present an extensive additional analysis of concept sensitivity (Section B) and Concept-Aware LoRA (Section C). We then demonstrate the generalizability of the proposed method using a general-domain dataset, Pascal VOC [12] (Section D). Additionally, we discuss class-aware dataset generation, which could be a promising future research direction for handling long-tailed distributions by leveraging text prompts (Section F). Finally, we provide additional quantitative and qualitative results (Section E) and detailed implementation details, including PyTorch-like pseudocode, to ensure reproducibility. Below is the table of contents for the supplementary material (Section G).

### v2

@@ 4 Experiments
In the following sections, we present the answers to the following research questions:
  1. 1.
    How does the generated dataset contribute to improving segmentors across various settings? (Section 4.2)
  2. 2.
    How important is image-domain alignment in enhancing segmentation performance? (Sections 4.2 and 4.3)
  3. 3.
    How does the T2I model’s behavior change when the target concept is modified? (Section 4.3)
  4. 4.
    How does concept-aware layer selection outperform hand-crafted baselines? (Section 4.4)
Figure 5: Qualitative comparison of image-label pairs. InstructPix2Pix only edits textures and thus cannot produce diverse scenes, while DATUM generates personalized but unlabeled images. DatasetDM suffers from viewpoint and style misalignment due to the lack of adaptation, and LoRA tends to memorize training examples. In contrast, CA-LoRA selectively learns style or viewpoint concepts, producing more diverse and better-aligned image-label pairs.
@@ 4.1 Experimental Setup
Implementation Details
We use SDXL [37] as the pretrained T2I model. For a fair comparison, we reimplement DatasetDM [51] with SDXL as well. We fix the rank of both the original LoRA and CA-LoRA at 64 and set the training iterations to 10k. We utilize the Style CA-LoRA for in-domain segmentation whereas Viewpoint CA-LoRA for domain generalization. We explore the selected proportion of CA-LoRA across 1%, 2%, 3%, 5%, and 10%. The diffusion timestep $t$ for identifying the desired concept has been searched in [1, 81, 201, 481] out of 1000 timesteps [56] and set to 81. Although our method requires additional CA-LoRA fine-tuning, it takes only about one hour on a single V100 GPU, adding negligible overhead compared to the 20-hour label generator training. The results of our hyperparameter search, along with details such as the label generator architecture, are provided in Appendices B, C and A.
Baselines
We compare our approach with prior dataset generation methods, including InstructPix2Pix [4], DATUM [3] [^1] For DATUM [3], we add one extra image per weather condition to meet the one-shot UDA setup and exclude it from the in-domain baseline, as it requires a UDA method., and DatasetDM [51]. For fine-tuning approaches, we further compare our concept-aware method with LoRA [20] and AdaLoRA [59] to demonstrate the benefit of selective fine-tuning. All fine-tuning methods are implemented on top of the DatasetDM framework.
@@
For baseline model, we train Mask2Former [5] on subsets of the Cityscapes [8] dataset at various fractions (0.3%, 1%, 3%, 10%, and 100%), which we denote as ‘Baseline’. Then, we generate 500 image-label pairs for all few-shot settings and use them as an additional dataset to fine-tune the baseline model, where we generate 3000 pairs for the fully-supervised setting. To prevent overfitting to the generated data, we mix real (e.g., samples from the Cityscapes dataset) and generated samples in equal proportions within each mini-batch.
@@
For the evaluation, we train on the Cityscapes [8] and test on ACDC [45], Dark Zurich (DZ) [44], BDD100K (BDD) [58], and Mapillary Vistas (MV) [33]. We improve the DG performance for urban-scene segmentation by building upon existing DG methods, including ColorAug, DAFormer [18] and HRDA [19]. To show the effectiveness of the generated dataset, we train a semantic segmentation model with each DG method on a combination of the 2975 Cityscapes image-label pairs and the 2500 generated image-label pairs from scratch. We generate 500 images for 5 adverse weather or lighting conditions including clear, foggy, night-time, rainy, and snowy.
@@ 4.2 Main Results on the Segmentation Benchmarks
Table 1 shows that the proposed CA-LoRA consistently outperforms all other methods across various data ratios. It improves mIoU by 2.30 at the 0.3% data fraction of Cityscapes, significantly surpassing DatasetDM, which yields only a 0.99 increase. Although the fine-tuning baselines (LoRA and AdaLoRA) show competitive performance at lower data fractions, they lag behind CA-LoRA at larger fractions (e.g., 10%, 100%) and struggle to generate informative samples beyond the source domain. As shown in Figure 5, InstructPix2Pix offers limited improvements because it mainly alters the texture of the scene and cannot meaningfully modify structural content. The consistent gains across data ratios highlight the robustness of CA-LoRA, demonstrating its effectiveness for urban-scene segmentation in both few-shot and fully supervised settings. We also observe consistent gains on PASCAL VOC [11], showing that CA-LoRA is applicable beyond urban-scene segmentation (see Appendix D).
@@
As shown in Table 2, the proposed method consistently outperforms existing segmentation dataset generation approaches across multiple DG methods [54, 18, 19]. Figure 5 further illustrates that CA-LoRA learns only viewpoint information from the source dataset (Cityscapes) while preserving the pretrained T2I model’s diverse style-generation capability. This results in strong generalization performance, particularly on challenging datasets such as ACDC and Dark Zurich, where adverse weather and lighting conditions drive the domain shift. In contrast, BDD100K and Mapillary Vistas differ mainly in scene mood or appearance, leading to milder shifts compared to ACDC and DZ. Finally, our approach improves performance not only over simple DG methods (ColorAug) but also when combined with advanced ones (DAFormer and HRDA), as our generated datasets provide semantically diverse content that label-to-image methods such as InstructPix2Pix and DGInStyle [22] struggle to reproduce.
4.3 Style-aware LoRA vs. Viewpoint-aware LoRA
This section provides an in-depth analysis of the CA-LoRA fine-tuned models (style and viewpoint), comparing them with the pretrained T2I model (0%, no fine-tuning; DatasetDM) and the LoRA fine-tuned model (100%, fine-tuning all layers; LoRA) used in the main results.
Figure 6: Comparison of Style CA-LoRA and Viewpoint CA-LoRA. Even when fine-tuning only 1% of the style-aware layers, image-domain alignment improves as quickly as fine-tuning 5–10% of the viewpoint-aware layers. In contrast, viewpoint CA-LoRA maintains strong text adherence even when fine-tuning 3–5% of its layers. Finally, style-aware layers show an advantage in image-label alignment, as their generative features exhibit a smaller domain gap.
@@
We evaluate the image-domain alignment of T2I models, which is crucial for generating an in-domain dataset. To measure this alignment, we adopt CMMD [21], as maximum mean discrepancy provides a reliable comparison for small image sets. Figure 6 (left) shows that the pretrained T2I model exhibits a large domain gap between real and generated images, whereas fine-tuning on Cityscapes effectively reduces this gap. Between the two CA-LoRA variants, Style CA-LoRA achieves comparable alignment while using only 10% of the parameters of the full LoRA. Moreover, fine-tuning just 1% of the style-aware layers yields image-domain alignment similar to fine-tuning 5–10% of the viewpoint-aware layers. Although the full LoRA (100%) achieves the best CMMD score, Figure 1 shows that it suffers from severe memorization, providing limited useful variation and resulting in only marginal segmentation improvements.
@@
Since generating diverse weather conditions (e.g., foggy, rainy) is crucial for improving DG performance, preserving the text adherence ability of the T2I model is essential. We measure the weather-conditional image generation performance by leveraging CLIP-Score [38], which can assess the similarity between the generated images and their input text prompts. Figure 6 shows the average CLIP-Score across four diverse weather conditions, based on 100 generated images per condition. The results indicate that the pretrained model achieves a high CLIP score for generating adverse weather conditions, whereas the Style CA-LoRA drastically fails to do so, as they primarily learn styles from the source dataset, which inherently includes its weather condition (e.g., clear-day). In contrast, the Viewpoint CA-LoRA effectively preserves the adverse weather conditional generation performance while learning the viewpoint from the source dataset.
@@
Generating reliable labels is a key requirement for constructing high-quality segmentation datasets. As shown in Figure 5, CA-LoRA achieves superior image–label alignment compared to the baselines by fine-tuning the T2I model for domain alignment. To quantitatively evaluate this alignment, we obtain pseudo–ground truth for the generated images using predictions from strong pretrained segmentors [5]. [^2] This is reasonable when the pretrained segmentors are trained on substantially larger datasets than the label generator. As shown in Figure 6, Style CA-LoRA achieves better image–label alignment than Viewpoint CA-LoRA, and increasing the number of fine-tuned layers further improves alignment. This improvement arises from the reduced domain gap between the training and inference-time generative features of the T2I model. However, increasing the number of fine-tuned layers is not always beneficial, as it can reduce scene diversity and produce less informative samples, ultimately yielding only marginal segmentation improvements.
Table 3: Ablation on parameter group selection for fine-tuning. The image domain alignment is measured by CMMD, and the final segmentation performance (mIoU, Cityscapes 0.3%).
Finetuning Group % Params. CMMD ($\downarrow$) Performance ($\uparrow$)
No Finetuning (DatasetDM) 0% 5.063 42.82
Q projections 25% 4.305 40.50 (-2.32)
K projections 25% 3.990 43.50 (+0.68)
V projections 25% 3.003 42.77 (-0.05)
OUT projections 25% 3.005 42.82 (+0.00)
All projections (LoRA) 100% 0.644 42.97 (+0.15)
Random selection 2% 0.783 43.24 (+0.42)
Concept-Awareness w/o norm. 2% 0.730 43.01 (+0.19)
Concept-Awareness (Ours) 2% 1.420 44.13 (+1.31)
4.4 Performance According to the Fine-tuning Parameter Group
To demonstrate the effectiveness of our parameter-selection approach, we conduct several ablation studies in the in-domain few-shot setting (Cityscapes 0.3%). We report improvements in both image-domain alignment and segmentation performance over the no-finetuning baseline (DatasetDM). Our evaluation includes hand-crafted projection–layer selection strategies (Q, K, V, OUT), random selection, and Concept-Awareness–based selection without diffusion-gradient normalization (see Equation 7).
As shown in Table 3, hand-crafted layer-selection strategies (Q, V, OUT) do not meaningfully improve either image-domain alignment or segmentation performance. Among them, only K projections show noticeable gains, yet they still fall short of our method. Fine-tuning all projection layers provides the best domain alignment, but it fails to generate informative samples due to memorization. Selecting a random subset of projection layers alleviates this issue to some extent but remains limited, underscoring the need for concept-aware selection. Concept-Awareness without normalization behaves similarly to full-layer fine-tuning, as it selects weights dominated by diffusion gradients; while parameter-efficient, it lacks true concept specificity. In contrast, our Concept-Awareness method achieves competitive domain alignment and yields the largest segmentation improvement (+1.31 mIoU over DatasetDM), demonstrating the effectiveness of concept-aware fine-tuning.
@@ 5 Conclusion and Future Work
This paper proposes Concept-Aware LoRA (CA-LoRA), a novel fine-tuning method that learns only the desired concepts (e.g., viewpoint or style) from the training dataset to generate segmentation datasets. By selectively identifying and updating only the weights relevant to the desired concepts, CA-LoRA enables the fine-tuned T2I model to produce well-aligned and informative samples. Moreover, it effectively leverages pretrained knowledge from T2I models to generate datasets, significantly improving segmentation performance across various settings, including in-domain (few-shot and fully-supervised) and domain generalization tasks. Furthermore, CA-LoRA excels at learning only the desired concepts from training data, even when it contains diverse information, including unnecessary concepts for dataset generation, making it applicable to various knowledge transfer and fine-tuning tasks.
Limitations
While we carefully construct synthetic segmentation datasets that achieve state-of-the-art performance among dataset generation approaches, the resulting performance gains may still appear marginal in certain settings. Nevertheless, adapting pretrained T2I models to better align with the target domain is crucial for dataset generation, and we believe our work provides a valuable foundation for principled approaches. Furthermore, since our method primarily focuses on incorporating style and viewpoint concepts that are informative for urban-scene segmentation, extending the concept-awareness framework to capture other types of concepts beyond style and viewpoint represents a promising direction for future research.
@@
In the supplementary material, we provide additional analyses, experiments, qualitative and quantitative results, and implementation details, including pseudocode, due to space limitations in the main manuscript. Specifically, we describe implementation details, including PyTorch-like pseudocode, to ensure reproducibility (Appendix A). Next, we illustrate an implicit bias of concept gradients (Appendix B) and additional experimental studies of our design choice of CA-LoRA (Appendix C). We then demonstrate the generalizability of the proposed method using a general-domain dataset, Pascal VOC [11] (Appendix D). Finally, we discuss class-aware dataset generation, which could be a promising future research direction for handling long-tailed distributions by leveraging text prompts (Appendix E). Below is the table of contents for the supplementary material.
Appendix A Implementation Details
A.1 Label Generator Architecture
Since DatasetDM is built on top of Stable Diffusion v1.5 [41], we simply adjust the feature dimensions in the projection layers to accommodate Stable Diffusion XL [37]. The detailed label generator architecture is illustrated in Figure 7. However, if all feature maps and cross-attention maps are used, the total number of channels increases significantly, leading to an unmanageable number of parameters in the projection layers during concatenation and projection. In summary, the feature maps are extracted from the last feature block at each resolution of the up-sampling blocks, while cross-attention maps are sampled at equal intervals (every 7 blocks) from the total 36 up-sampling blocks (i.e., 1st, 8th, … 29th, 36th).
Furthermore, as shown in Figure 7, Stable Diffusion XL has only three resolution levels, compared to four resolution levels of the Stable Diffusion v1.5 architecture in the original DatasetDM. In the Mask2Former structure, feature maps from the pixel decoder, excluding the largest resolution, are fed into the transformer decoder. While the original design used three-resolution feature maps, only two were utilized in this case. Thus, while DatasetDM provides three-resolution feature maps three times for 9 transformer decoder blocks, we provide two-resolution feature maps five times, leading to a total of 10 transformer decoder blocks. Importantly, to ensure a fair comparison, the reported scores for DatasetDM were obtained using a re-implemented version based on SDXL with the same modifications.
A.2 Detailed Architecture of CA-LoRA
Overview (Figure 8 (a))
The fundamental group of weights used to measure concept awareness is the linear projection layer. We selectively adapt the pretrained weights layer by layer within the projection layers. Practically, the projection layers in the multi-head attention block are integrated across heads, we split them to structurally distinguish the weights for each projection. Consequently, we attach LoRA layers projection-wise, as shown in Figure 8.
Projection-wise CA-LoRA (Figure 8 (b))
$\hskip-28.45274pt\Delta W_{\text{OUT}}=B\begin{bmatrix}&&&\\ A_{1}&A_{2}&\cdots&A_{h}\\ &&&\end{bmatrix}$, (8)
The projection-wise LoRA is represented in Figure 8 and Algorithm 4.
A.3 Hyperparameters and Pseudocode
Hyperparameters (Tables 12, LABEL:, 13, LABEL:, 14, LABEL: and 15)
Table 4: Hyperparameter search of layer proportions. Image-domain alignment (CMMD) and final segmentation performance.
Proportion CMMD Performance (mIoU)
0% (Pretrained) 5.063 42.82
1% 1.618 43.77
2% 1.420 44.13
3% 1.021 43.94
5% 1.105 43.81
10% 0.686 43.36
100% (Original LoRA) 0.644 42.97
Pseudocode
We also provide PyTorch-like pseudocode [36] for key algorithms to effectively support reproducibility. Concept awareness (Algorithms 1 and 2) The concept awareness algorithm is demonstrated in Algorithm 1. Conducting concept awareness requires several helper functions, as shown in Algorithm 2. CA-LoRA (Algorithms 3 and 4) The CA-LoRA algorithm is divided into two parts: the forward function and the declaration function. The forward pass of CA-LoRA is presented in Algorithm 3, while the declaration function, along with the selected layers, is illustrated in Algorithm 4.
Figure 9: Visualizing the gradients and concept awareness. (a) Implicit gradient biases across network layers. (b) Concept awareness, computed by normalizing each concept gradient with respect to the original diffusion-loss gradient. Style-aware and viewpoint-aware weights are highlighted in red and blue, respectively. Although raw concept gradients alone cannot clearly separate different concepts, normalization with diffusion gradients yields a more distinguishable concept-awareness signal.
Appendix B Visualizing Concept Awareness
B.1 The Implicit Bias of Gradients Across Layers
Observation (Figure 9 (a))
We compute the awareness scores for each layer by using the norm of the gradient. However, the gradient norm cannot be uniformly scaled across different head types (Q, K, V, OUT), attention types (self, cross), and layers (shallow, deep). To address this, we construct a base gradient to scale the concept loss gradient by referencing the gradient of the original diffusion loss, as described in Section 3.2. We visualize the gradients of both the concept losses and the original diffusion loss in Figure 9. In this visualization, we separate the self-attention and cross-attention layers to provide clearer distinctions, which differs from the approach in the main paper.
Normalizing Gradients (Figure 9 (b))
Therefore, we normalize the gradients using the gradients calculated from the original diffusion loss, as discussed in Section 3.2 and shown in Figure 3. As shown in Figure 9 (a), the gradients calculated from the style concept loss and viewpoint concept loss are similar. However, the gradient increase ratio can differ significantly, as illustrated in Figure 9 (b).
Figure 10: Measured concept awareness according to the various prompt augmentation. The highlighted concept-aware layers for each concept (style and viewpoint) remained largely consistent regardless of the prompt augmentation, demonstrating the robustness of concept awareness to variations in prompt design.
Figure 11: Visualizing concept awareness across different noise timesteps. It shows that the 81st timestep stands out with a significantly distinct concept awareness score between style and viewpoint awareness compared to the other timesteps.
Figure 12: Qualitative comparison across different noise timesteps. According to the various noise timestep, 81st timestep represents the best concept awareness, qualitatively. The style of the generated images by Style CA-LoRA is well-aligned, while the generated images by Viewpoint CA-LoRA contain diverse styles.
B.2 Concept Awareness per Prompt Augmentation
We conduct an additional analysis of the robustness of defining desired concepts based on prompt augmentation. As demonstrated in Section 3.2, we select several hand-crafted prompt augmentations tailored to the desired concept. While this method is flexible and can be generalized to different concepts, it may introduce instability in prompt augmentation. To assess its robustness, we measure awareness across various prompt augmentations. Specifically, we provide five prompt augmentations for each desired concept (style, viewpoint) derived from the original prompts, as shown below.